# Lean classical-quantum hybrid neural network model for image classification

### v2

@@ 3.3 Visual Explanations
Neural network models are frequently considered as black-box systems due to their inherently complex nonlinear transformations and opaque internal structural parameters. This lack of transparency prompts researchers to investigate the focal points of these models. To enhance our understanding of the model’s focus in this paper, we introduce Gradient-weighted Class Activation Mapping (Grad-CAM), an interpretability technique that leverages gradient information to identify regions in an image that contribute most significantly to the model’s predictions for a specific class. By combining gradients with activation maps from convolutional layers to generate Class Activation Maps (CAMs), areas with substantial activation are superimposed onto the original image, thereby highlighting regions that play a critical role in class prediction. We visualize the training process through heatmaps, which provide clearer insights into shifts in the model’s attention direction. The detailed computational formula is presented subsequently[54].
@@
The visualization interpretation results of the LCQHNN on the MNIST dataset are depicted in Figure 6. We selected the last convolutional layer as the target layer to observe the significant areas the model uses for predictive concepts. Figure 6(a) presents the input original image, while Figures 6(b), 6(c), and 6(d) illustrate the visualization interpretation results of the model at different training stages. We use the heatmaps to represent the importance of various regions in the image for the target category. The closer the heatmap pixel color is to dark red, the more critical the model considers that area for predicting the target category. Initially, for the untrained model (Figure 6(b)), the visualization shows that the model’s focus is scattered and chaotic, with no discernible pattern. This indicates that the model’s initial cognition of the target category is inaccurate, and it has not yet learned effective features related to prediction. Subsequently, the model in the middle of training (Figure 6(c)) begins to learn some feature information from the training set data, the model starts to focus on the edge contours of the handwritten digits, signifying that it is gradually understanding some important features of the digit shapes through training. Finally, the model after completion of training (Figure 6(d)) shows that the areas of focus largely coincide with the shape of the handwritten digits. This demonstrates that the model has learned the feature information of the data and associated these features with the target categories after training. The model’s ability to accurately identify and locate the important areas of the handwritten digits confirms the effectiveness of the LCQHNN on this task. Moreover, this indicates the practicality of the VQCs in neural networks, enhancing the model’s expressive power and performance by combining classical and quantum computing advantages. By leveraging the unique features of quantum computing, VQCs can offer advantages in handling complex data patterns and tasks, providing superior predictive and interpretive capabilities.
@@ 4 Conclusion
In this study, we propose an innovative hybrid model that integrates classical and quantum computing, named LCQHNN, and conduct an in-depth comparative analysis of it. The novelty of the LCQHNN model lies in its combination of the powerful feature extraction capabilities of traditional CNNs with the classification potential of VQCs. Specifically, the model first utilizes CNNs to extract feature information from images and then effectively hands off these features to VQCs for precise classification tasks. Experimental results demonstrate that the improved classical-quantum hybrid model exhibits superior performance in classification tasks across multiple public datasets, with enhancements in both recognition accuracy and convergence speed compared to traditional CNNs models with the same parameters. In the binary classification task of MNIST involving digits 0 and 1, the model achieved 100% accuracy at the fastest speed for image recognition. On the FashionMNIST dataset, LCQHNN surpassed CNNs in both convergence rate and final accuracy, including those with a richer set of parameters. These results indicate that the LCQHNN model, which combines classical networks and quantum computing, can achieve higher precision and provides a reliable, flexible, and scalable deep learning approach for image classification tasks. Additionally, we observed the model’s visual interpretation results, which show that the model effectively captured key data features during the training process and established associations with corresponding categories. This visual analysis highlights the capability of LCQHNN to enhance performance by fine-tuning its learning trajectory when dealing with complex image data. Furthermore, it provides compelling evidence to explore the emerging potential of quantum computing in deep learning, paving the way for innovative and more accurate hybrid quantum-classical models.
@@ Acknowledgments
Code availability: The code for this study can be obtained from the corresponding author upon request.

### v3

@@ 3.3 Visual Explanations
Neural network models are frequently considered as black-box systems due to their inherently complex nonlinear transformations and opaque internal structural parameters. This lack of transparency prompts researchers to investigate the focal points of these models. To enhance our understanding of the model’s focus in this paper, we introduce Gradient-weighted Class Activation Mapping (Grad-CAM), an interpretability technique that leverages gradient information to identify regions in an image that contribute most significantly to the model’s predictions for a specific class. By combining gradients with activation maps from convolutional layers to generate Class Activation Maps (CAMs), areas with substantial activation are superimposed onto the original image, thereby highlighting regions that play a critical role in class prediction. We visualized the training process of MNIST and FashionMNIST through heat maps, which provide clearer insights into shifts in the model’s attention direction. We did not experiment on CIFAR-10 because the heat map visualization technique we used is specifically designed for grayscale images with a consistent pixel format. CIFAR-10 consists of color images with a different spatial resolution ($32\times 32$ pixels) and channel structure (RGB), which makes it incompatible with our current visualization framework. The detailed computational formula is presented subsequently[56].
@@
The visualization interpretation results of the LCQHNN on the MNIST dataset are depicted in 7. We selected the last convolutional layer as the target layer to observe the significant areas the model uses for predictive concepts. Figure 7(a) presents the input original image, while Figures 6(b), 6(c), and 6(d) illustrate the visualization interpretation results of the model at different training stages. We use the heatmaps to represent the importance of various regions in the image for the target category. The closer the heatmap pixel color is to dark red, the more critical the model considers that area for predicting the target category. Initially, for the untrained model (Figure 7(b)), the visualization shows that the model’s focus is scattered and chaotic, with no discernible pattern. This indicates that the model’s initial cognition of the target category is inaccurate, and it has not yet learned effective features related to prediction. Subsequently, the model in the middle of training (Figure 7(c)) begins to learn some feature information from the training set data, the model starts to focus on the edge contours of the handwritten digits, signifying that it is gradually understanding some important features of the digit shapes through training. Finally, the model after completion of training (Figure 7(d)) shows that the areas of focus largely coincide with the shape of the handwritten digits. This demonstrates that the model has learned the feature information of the data and associated these features with the target categories after training. The model’s ability to accurately identify and locate the important areas of the handwritten digits confirms the effectiveness of the LCQHNN on this task. Moreover, this indicates the practicality of the VQCs in neural networks, enhancing the model’s expressive power and performance by combining classical and quantum computing advantages. By leveraging the unique features of quantum computing, VQCs can offer advantages in handling complex data patterns and tasks, providing superior predictive and interpretive capabilities.
@@ 4 Conclusion
In this study, we propose an innovative hybrid model that integrates classical and quantum computing, named LCQHNN, and conduct an in-depth comparative analysis of it. The novelty of the LCQHNN model lies in its combination of the powerful feature extraction capabilities of traditional CNNs with the classification potential of VQCs. Specifically, the model first utilizes CNNs to extract feature information from images and then effectively hands off these features to VQCs for precise classification tasks. Experimental results demonstrate that the improved classical-quantum hybrid model exhibits superior performance in classification tasks across multiple public datasets, with enhancements in both recognition accuracy and convergence speed compared to traditional CNNs models with the same parameters. This means that the LCQHNN model combining classical networks and quantum computing can achieve higher accuracy and provide a reliable, flexible, and scalable deep learning method for image classification tasks. Additionally, we observed the model’s visual interpretation results, which show that the model effectively captured key data features during the training process and established associations with corresponding categories. This visual analysis highlights the capability of LCQHNN to enhance performance by fine-tuning its learning trajectory when dealing with complex image data. The LCQHNN framework holds significant promise for practical deployment in scenarios requiring high-precision and efficient image analysis. For instance, in medical diagnostics, it could enhance the detection of subtle anomalies in radiological images (e.g., early-stage tumors in MRI scans) by leveraging quantum-enhanced classification to reduce false negatives. In autonomous systems, such as self-driving vehicles, LCQHNN’s compactness and accuracy could improve real-time object recognition under dynamic environmental conditions. Despite its advantages, practical adoption of LCQHNN faces several challenges. First, current quantum hardware limitations, such as qubit decoherence, gate fidelity, and limited qubit counts, constrain the scalability of VQCs for large-scale datasets. Second, The LCQHNN’s reliance on transferring classical features (extracted by CNNs) to quantum circuits introduces a critical bottleneck: the dimensionality mismatch between high-dimensional classical features and limited qubit resources. Current quantum encoders struggle to efficiently map complex image features (e.g., 512-dimensional CNN outputs) into low-qubit quantum states without significant information loss. This trade-off between encoding fidelity and quantum resource overhead may degrade performance in real-world scenarios with high-resolution inputs. As quantum hardware advances, LCQHNN could emerge as a transitional framework that harmonizes classical and quantum paradigms, advancing AI solutions in domains where the interplay of precision and efficiency is critical.
@@ Acknowledgments
Code availability: The code and implementation details for this study are publicly available at https://github.com/LAAAAAAAAA/LCQHNN.